OpenStudy (anonymous):

Ratio question : The ratio of the ages of Anna and Emma is 3:5 and the ratio of the ages of Emma and Nicholas is 3:5 . What is the ratio of Anna's age to Nicholas age? I have no idea how to deal with this type of question.

OpenStudy (anonymous):

Anna age = A Emma age = E Nicholas age = N If Anna/Emma = 3/5 , you can say that Anna age is 3K and Emma age is 5K (K a constant). Ok, Emma / Nicholas = 3/5, change Emma to 5K and Nicholas to N. 5K/N = 3/5 25K = 3N N = 25K/3 You want the ratio of Anna and Nicholas. If Anna = 3K and Nicholas = 25K/3, the ratio is : 3K / (25K/3) = 9/25

OpenStudy (anonymous):

Thank you!, that was a great answer , I still have trouble getting the logic behind it . I will have to do more of these problems

OpenStudy (anonymous):

There are others ways to solve this, you can solve without using the constant k, but for me it's more difficulty. Wanna see that?

OpenStudy (anonymous):

Again, Anna age = A Emma age = E Nicholas age = N \[\frac{ A }{ E } = \frac{ 3 }{ 5 }\]\[5A=3E\]\[E = \frac{ 5A }{ 3 }\]. Ok, and \[\frac{ E }{ N } = \frac{ 3 }{ 5 }\] \[5E=3N\] Change \[E \to \frac{ 5A }{ 3 }\] You will have: \[5*(\frac{ 5A }{ 3 }) = 3N\]\[\frac{ 25A }{ 3 }=3N\]\[25A=9N\]\[\frac{ A }{ N }=\frac{ 9 }{ 25 }\]

OpenStudy (anonymous):

oh wow , I actually got this explanation better. Thanks a lot!!!